OpenStudy (elite):

Simplify |-3| + 6 ÷ (4 - 1) • 8.

OpenStudy (wendy.ivette11714):

\[|−3|+\frac{ 6 }{ 4-1 } \times 8\] \[= 3 + \frac{ 6 }{ 3 }(8)\] \[=3+(2)(8)\] \[=3+16\]

OpenStudy (devonhoward15):

@Wendy.Ivette11714 i only gave one anwser

OpenStudy (wendy.ivette11714):

What is 3+16 @elite

OpenStudy (elite):

ok thanks

OpenStudy (elite):

its 19

OpenStudy (wendy.ivette11714):

Yes, 19 is the answer ^_^

OpenStudy (elite):

thanks

OpenStudy (elite):

your good at this

OpenStudy (wendy.ivette11714):

your welcome

OpenStudy (wendy.ivette11714):

thanks

OpenStudy (elite):

your names wendy